OpenStudy (anonymous):

The average of 6 integers is less than 30. If the average of the first 5 integers is 25, what is the greatest possible value of the 6th integer?

OpenStudy (anonymous):

Sounds like an SAT question.. :-p Do you know the formula for finding an average?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

add up all the numbers then divide by how many u added

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

also known as the mean :D

OpenStudy (anonymous):

@gaara438125 , exactly, so you can start with the first statement and say (using the letters a-f as the six integers) \[\large \frac{a+b+c+d+e+f}{6}<30.\]

OpenStudy (anonymous):

ya i know mean just not a formula

OpenStudy (anonymous):

(by 'formula' I really just meant, do you know how to do it, and how to write out the steps)

OpenStudy (anonymous):

Now, the second statement says \[\large \frac{a+b+c+d+e}{5}=25.\] See where to go from here, @evergirl ?

OpenStudy (anonymous):

I got 54

OpenStudy (anonymous):

That seems to check out. Good job!